At age 40, Nate purchased an annuity from an insurance company for $161,280. According to the terms of the contract, Nate will receive $2,000 per month for life beginning at age 66. Show all calculations using good form.

                  Required:

  1. Assume that Nate did not make any withdrawals before he reached age 66. Nate begins receiving monthly annuity payments on 6/1/2020, the date he turns age 66. How much gross income does Nate recognize in 2020?

                  2.    Assume that Nate lives 23 years after the annuity start date. How much gross income does Nate recognize in the 21st year?

                  3.    Assume that Nate turned age 66 on 6/1/2010. Further, assume that Nate dies on 5/31/2020 after collecting monthly annuity payments for exactly 10 years. As a result of the annuity contract, how much is Nate’s deduction on his 2020 income tax return? How much is included in gross income on Nate’s 2020 income tax return?
Answer #1

Answer

Part1.

Gross Income for 2020

= 7*2000 = 14,000

Part 2. Gross Income in 21st year = 12*2000 = 24,000

Part 3. Gross Income for 2020 = 5*2000 = 10,000

Deductions = Nil

Deductions: Qualified annuities are used in connection with tax-advantaged retirement plans, such as 401(k) plans, Section 403(b) retirement plans (TSAs), or IRAs. Contributions to nonqualified annuities are made with after-tax dollars--premiums are not deductible from gross income for income tax purposes.